# bibliography.bib

### v2

@@ Abstract
Multilevel sampling methods, such as multilevel and multifidelity Monte Carlo, multilevel stochastic collocation, or delayed acceptance Markov chain Monte Carlo, have become standard uncertainty quantification tools for a wide class of forward and inverse problems. The underlying idea is to achieve faster convergence by leveraging a hierarchy of models, such as partial differential equation (PDE) or stochastic differential equation (SDE) discretisations with increasing accuracy. By optimally redistributing work among the levels, multilevel methods can achieve significant performance improvement compared to single level methods working with one high-fidelity model. Intuitively, approximate solutions on coarser levels can tolerate large computational error without affecting the overall accuracy. We show how this can be used in high-performance computing applications to obtain a significant performance gain.
As a use case, we analyse the computational error in the standard multilevel Monte Carlo method and formulate an adaptive algorithm which determines a minimum required computational accuracy on each level of discretisation. We show two examples of how the inexactness can be converted into actual gains using an elliptic PDE with lognormal random coefficients. Using a low precision sparse direct solver combined with iterative refinement results in a simulated gain in memory references of up to 3.5×3.5$\times$3.5 × compared to the reference double precision solver; while using a MINRES iterative solver, a practical speedup of up to 1.5×1.5$\times$1.5 × in terms of FLOPs is achieved. These results provide a step in the direction of energy-aware scientific computing, with significant potential for energy savings.
@@ 1 Introduction
Suppose we are interested in sampling from a probability distribution of a certain quantity $Q$, which depends on the (infinite-dimensional) solution of a partial differential equation (PDE) or a stochastic differential equation (SDE). In most cases, direct access to $Q$ is unavailable. Instead, a numerical model is used to obtain a finite-dimensional approximation $Q_{L}$ of the quantity $Q$. With increasing $L$ the accuracy of the model increases, but so does the cost of computing the solution. This can be a finite element or finite difference method in the case of PDEs, or an Euler-Maruyama discretisation for SDEs. A first idea to approximate the unavailable distribution of $Q$ is to choose a high-fidelity model $Q_{L}$ with $L$ sufficiently large and to use this model for sampling. Instead, the key idea of multilevel and multifidelity sampling methods, such as multilevel Monte Carlo (MLMC) [giles2015multilevel], multilevel stochastic collocation [teckentrup2015multilevel], or delayed acceptance MCMC  [christen2005markov], is to use a hierarchy of models $Q_{0},\ldots,Q_{L}$ with increasing accuracy. Then, by combining all the samples from models $Q_{0},\ldots,Q_{L}$ in a suitable way, a significant cost gain can be achieved.
Convergence of the aforementioned multilevel schemes has been thoroughly studied in various settings. The analysis is typically concerned with balancing the sampling and bias error to achieve an optimal performance. This paper adds one additional layer to the analysis, which has, to our knowledge, largely been neglected. In practice, the finite-dimensional model $Q_{l}$ is never solved exactly, rather an approximation $\tilde{Q}_{l}$ is obtained on a computer. We show that multilevel methods often admit safe use of inexact computations on coarse levels $l$ without affecting overall sampling accuracy. An elegant way to achieve the given low accuracy of the inexact solution $\tilde{Q}_{l}$ is the use of iterative procedures, such as Krylov subspace methods or iterative refinement. Efficient implementation of these techniques may lead to significant gains in terms of both time and energy in high performance computing applications [haidar2018design, haidar2018harnessing, carson2015communication].
To give a concrete example, we consider the MLMC method and rigorously show how the required computational accuracy is tied to the model accuracy at level $l$. We further establish a theoretical framework for quantifying the computational error by choosing a suitable error model. To fully leverage it, one needs to determine where the dominant computational cost lies. In many models the dominant cost lies in the solution of linear algebraic equations, which is why, for our case study, we choose an uncertainty quantification problem given by an elliptic PDE with lognormal random coefficients. Another example, recently studied in [giles2024rounding], addresses the cost of generating random variables, which can also be significant in some cases.
With this motivation in mind, we will consider a model problem of the form
$-\nabla\cdot\bigl{(}a(\cdot,\omega)\nabla u(\cdot,\omega)\bigl{)}=f(\cdot,\omega).$
An elliptic PDE of this form depending on a random parameter $\omega$ arises, for example, in uncertainty quantification of groundwater flow [cliffe2011multilevel]. We are interested in estimating the value of a quantity of interest (QoI), given as a function $G$ of the solution $u$, such as the expected value of a given functional of the solution, e.g., the expected value of the value of the solution at a certain point in the domain. The coefficient $a$ and the right-hand side $f$ are assumed to be (infinite-dimensional) random fields. In practice, they also have to be approximated and depend on a high-dimensional random parameter vector $\omega$. Due to the high-dimensionality, the numerical approximation of the quantity of interest becomes exponentially difficult, which is known as the curse of dimensionality; see, e.g., [giles2015multilevel, nobile2008sparse].
Thus we are interested in sampling from the unavailable distribution of the quantity $Q=G(\omega)$ to compute statistics of this distribution, e.g., the mean $\mathbb{E}[Q]$. To obtain a computable approximation $Q_{L}$ of $Q$ we choose the finite element method (FEM). This introduces in our approximation the so-called discretisation error, referred to also as model error or bias. After the PDE has been discretised, we have to face the challenge of approximating $\mathbb{E}[Q_{L}]$. To this end, Monte Carlo (MC) methods are often used, which introduces a sampling error. In the standard Monte Carlo method we need to take a sufficient number of samples of the discrete solution computed on a sufficiently fine mesh in order to bound both the sampling and the discretisation error. A significant variance reduction can be achieved if the samples are taken on a hierarchy of discretisation levels. This is the underlying idea of the MLMC method [heinrich2001multilevel, cliffe2011multilevel, giles2015multilevel]. The MLMC method dramatically reduces the variance of the resulting estimator while preserving the overall cost and it allows us to efficiently balance the two errors — discretisation error and sampling error.
In our model problem, four main numerical procedures have to be performed. The first is sampling the random fields $a(\cdot,\omega)$ and $f(\cdot,\omega)$, $\omega\in\Omega$; the second is assembling the FE stiffness matrix; the third corresponds to solving the associated linear system; and the final one involves computing the MLMC estimate itself. In our case, the cost of sampling random variables is small, and so is the cost of computing the MLMC estimate itself. Moreover, using inexact computations to compute the MLMC estimate could result in violating the MLMC telescopic property, so it is not of great interest. The stiffness matrix assembly cost may be significant, but it can be implemented in linear complexity (see, for example, [weinberg2018fast]), so that it is not asymptotically dominant. Nevertheless, efficient implementation of the stiffness matrix assembly may also be of interest and bring a significant reduction in solution time. However, in this work, we focus on the dominant cost of solving the resulting linear system and how to reduce it using inexact computations. The error introduced by solving the linear system will be referred to as the computational error. It stems from the fact that the system is solved with an inexact linear solver, yielding an approximate solution $\tilde{Q}_{l}$ of the model $Q_{l}$, as discussed above. To balance the computational error with the sampling and discretisation errors, we determine the minimum computational accuracy required for solving the linear system, which is a general solver-independent result. In practice, how inexact computations can be exploited depends on the particular solver.
To demonstrate the flexibility of our approach, we consider both direct and iterative solvers and discuss how to exploit inexactness in each of them. For iterative solvers, we suggest to determine the optimal number of iterations via a suitable stopping criterion. For direct solvers, we use mixed precision arithmetic and employ the technique of iterative refinement; cf. [higham2002accuracy, vieuble2022mixed]. In recent years, there has been growing interest in the selective use of low precision floating point arithmetic to accelerate scientific computations while maintaining acceptable levels of accuracy; cf. [higham2022mixed].
There are a couple of works that go in the direction of this manuscript. In [brugger2014mixed], the authors explore the use of lower precision on field programmable gate arrays (FPGA) in the MLMC method, using a stochastic differential equation (SDE) discretised by the Euler method as the model problem. The recent paper by Giles and Sheridan-Methven [giles2024rounding] is concerned with the analysis of rounding errors in generating random variables in the context of SDEs, again with application to multilevel Monte Carlo. We approach this topic from a different perspective. Namely, we
@@
  • propose a novel adaptive algorithm, determining the minimum required computational accuracy on each level of discretisation using a-priori error estimates with no additional cost (Section 4.4);
  • provide a theoretical basis for applying the adaptive algorithm to the model elliptic PDE with random coefficients and a random right-hand side (Section 4.3);
  • demonstrate the efficiency of the adaptive algorithm through a sequence of numerical examples, achieving up to a factor of about 3.5×3.5$\times$3.5 × in simulated memory gain with iterative refinement and a speedup by a factor of about 1.5×1.5$\times$1.5 × in terms of floating point operations with an iterative solver. A cost analysis and possible use in energy-efficient scientific computing are presented in Section 4.5.
The manuscript is divided into five sections. In the second section, we discuss linear solvers and various types of errors induced by their use. We give an overview of floating point arithmetic, the technique of iterative refinement, as well as some convergence results of Krylov subspace methods. The third section introduces the elliptic model problem, as well as its numerical solution by finite elements. Selected FE convergence results are presented and the convergence of the FEM using inexact solvers is discussed. In the fourth section we give a brief overview of the standard MLMC method, and the novel analysis of the computational error in MLMC along with the adaptive algorithm and the cost analysis. Finally, the numerical results along with a thorough discussion are given in Section 5.
2 Linear solvers and inexact computations
@@ 2.1 Floating point arithmetic
A floating point (FP) number system $\mathbb{F}$ is a finite subset of real numbers whose elements can be written in a specific form; see [higham2002accuracy] for a thorough description. All computations in FP arithmetic are assumed to be carried out under the following standard model: For all $x,y\in\mathbb{F}$
@@
where $\varepsilon$ is the unit roundoff. This assumption allows us to analyse the error of a given algorithm. For simplicity, we will often abbreviate “the floating point arithmetic format with unit roundoff $\varepsilon$” to “the precision $\varepsilon$”.
The standard model is valid in particular for IEEE arithmetic, a technical standard of floating point arithmetic, which assumes the preliminaries above and adds other technical assumptions; see [higham2002accuracy] for an overview. The IEEE standard defines several basic formats. In this work, we use formats both standardized and not standardised by IEEE. All of the formats we will use are hardware-supported, namely by the NVIDIA H100 SXM GPU (for specifications see [nvidia_h100_specs]). To be specific, in this manuscript we use quarter (q43), half, single, and double precision. Half, single, and double are basic IEEE formats. The quarter precision format we use has $4$ exponent bits and $3$ significand bits, which means storing one number requires $8$ bits including the sign. The unit roundoff of quarter precision is $2^{-4}$ and its range is $10^{\pm 2}$.
@@ 2.2 Krylov subspace methods
Krylov subspace methods are a powerful class of iterative methods used for solving large-scale linear systems of equations and eigenvalue problems, particularly those involving sparse or structured matrices. Our application yields matrices which are symmetric positive definite (see Section 3), which makes the conjugate gradient (CG) and MINRES methods two suitable Krylov subspace methods for our purpose. The following overview of the methods is adapted from [greenbaum1997iterative, Section 3.1].
Consider a linear system $Ax=b$ with symmetric positive definite $A\in\mathbb{R}^{n\times n}$. Let $x_{0}\in\mathbb{R}^{n}$ be an inital estimate of the solution, $e_{0}=x-x_{0}$ the initial error, and $r_{0}=b-Ax_{0}$ the initial residual. Then, in each iteration $k$, CG minimizes the $A$-norm of the error and MINRES minimizes the Euclidean norm of the residual over the Krylov subspace $x_{0}+\text{span}\{r_{0},A^{2}r_{0},\dots,A^{k-1}r_{0}\}.$
The stopping criterion we set is given by the relative residual and is motivated by our specific use of the linear solver; see Sections 4.2 and 5. To be precise, we require the solution produced by the iterative algorithm to to satisfy (1) for a given $\delta_{e}>0$. This motivates our choice to use the MINRES method in this work. It has been shown that residuals produced by the CG and MINRES methods on Hermitian matrices are closely related (see [greenbaum1997iterative, Exercise 5.1]), and therefore both methods are expected to perform well for our model problem to be introduced in Section 3. Using the stopping criterion given by the relative residual will allow us to apply abstract error estimates from Section 4.2 to MINRES.
It is clear that in exact arithmetic, MINRES converges to the true solution in a finite number of iterations. In finite precision arithmetic, this is not the case, and the convergence analysis is nontrivial; see for example [greenbaum1997iterative, Chapter 4]. We therefore perform all calculations in the MINRES method in double precision. In order to apply the abstract error estimates from Section 4.2 to MINRES, we assume $\widehat{x}_{k}\approx x_{k}$, where $\widehat{x}_{k}$ and $x_{k}$ are the MINRES solutions from $k$-th iteration computed in floating point and exact arithmetic, respectively. An alternative approach to performing all calculations in high precision would be to employ a Krylov subspace method coupled with iterative refinement; see Section 2.3 for an overview of iterative refinement. Iterative refinement as a technique to improve accuracy of a Krylov subspace method was used, for example, in [carson2017new] and [carson2018accelerating] with the GMRES method.
@@ 2.3 Iterative refinement
Iterative refinement is a technique used to enhance the accuracy of a numerical solution to a linear system by computing a residual vector and then applying a correction to the current approximation to reduce the error; see [higham2002accuracy, Section 12] for a detailed overview. In this way, the process can be repeated iteratively until the limiting level of accuracy is achieved. The particular version of iterative refinement which we use in this work is adapted from [carson2018accelerating]. Consider a linear system $Ax=b$ where $A\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$. The iterative refinement algorithm works explicitly with the following three precisions:
  • $\varepsilon$ is the (working) precision at which the data $A$, $b$ and solution $x$ are stored,
  • $\varepsilon_{f}$ is the precision at which the factorisation of $A$ is computed,
  • $\varepsilon_{r}$ is the precision at which residuals are computed.
Implicitly, the algorithm contains a fourth precision, $\varepsilon_{s}$, which is the precision at which the correction equation is effectively solved. Note that in this subsection the symbol $\varepsilon_{s}$ does not denote single precision and we use it here to keep the indices in accordance with [carson2018accelerating], where the precision are denoted by $u_{r}$, $u$, $u_{f}$, and $u_{s}$, respectively.
All of the precisions $\varepsilon_{r}$, $\varepsilon$, $\varepsilon_{s}$, and $\varepsilon_{f}$ will take, in the cases of interest, one of the values of quarter, half, single, or double precision, and it is assumed $\varepsilon_{r}\leq\varepsilon\leq\varepsilon_{s}\leq\varepsilon_{f}$. We present Algorithm 2.1 which is a special case of [carson2018accelerating, Algorithm 1.1]. As mentioned above, $\varepsilon_{s}$ is the precision at which the equation in step 8 is solved. In our case, $\varepsilon_{s}=\varepsilon_{f}$, since the factorisation is carried out in precision $\varepsilon_{f}$ and we always consider the factorisation method to be the Cholesky decomposition.
@@
0:  $A\in\mathbb{R}^{n\times n}$, $b\in\mathbb{R}^{n}$, both in precision $\varepsilon$, tolerance $\delta_{e}>0$.
@@
1:  Factorise $A$ in precision $\varepsilon_{f}$.
2:  Solve $Ax_{0}=b$ in precision $\varepsilon_{f}$ by substitution and store $x_{0}$ at precision $\varepsilon$.
@@
4:     Compute $r_{i}=b-Ax_{i}$ at precision $\varepsilon_{r}$ and round $r_{i}$ to precision $\varepsilon_{s}$.
@@
8:     Solve $Ad_{i}=r_{i}$ by (forward/backward) substitution at precision $\varepsilon_{f}$ or $\varepsilon$ using the factorisation computed in step 1 and store $d_{i}$ at precision $\varepsilon$.
9:     $x_{i+1}=x_{i}+d_{i}$ at precision $\varepsilon$.
@@
The precision to be used in step 8 of the algorithm is always specified when iterative refinement is used. By using a precision higher than the factorisation precision in step 8 of the algorithm, i.e., $\varepsilon_{s}<\varepsilon_{f}$, no numerical benefit from the perspective of the limiting tolerance for the residual norm $\|{r_{i}}\|/\|{b}\|$ can be obtained. However, it turns out that using higher precision in step 8 can help us in practice to reach the same desired tolerance faster with a small additional cost. In our case, there is also no reason to use extra precision to compute the residual in step 4 of the algorithm, since we are only interested in bounding the backward error. Computing the residual in higher precision is useful if bounding the forward error is of interest; see [carson2018accelerating].
@@
In this work we use Algorithm 2.1 extensively to achieve the desired accuracy of the computed solution. We obtain additional benefits from using iterative refinement in terms of the cost. The point is to reuse the factorisation computed in step 1 of the algorithm in step 8 of the algorithm. If we reuse the factorisation as suggested, we can improve the accuracy of the solution with a small additional cost, because the factorisation is typically the dominant part in terms of the required number of operations; see [vieuble2022mixed, Section 2.2].
@@
The stopping criterion in Algorithm 2.1, similarly as in Krylov subspace methods, is given by the relative residual norm. It is again motivated by our specific use of iterative refinement; see Sections 4.2 and 5. From the definition of the stopping criterion in step 5 it follows that if Algorithm 2.1 converges, then the produced solution is computed effectively to precision $\delta_{e}$ in the sense of (1). This will allow us to apply abstract error estimates from Section 4.2 to iterative refinement. The convergence of Algorithm 2.1 is analysed thoroughly in [carson2018accelerating].
The convergence properties of iterative refinement for linear systems stemming from finite element discretisations of elliptic PDEs will be investigated in Section 3.2.
2.4 Energy savings through iterative refinement
We finish this section by discussing how using low precision can translate into energy or runtime savings. In the past decades, power has become the principal constraint of computing performance [horowitz20141]. On the hardware level, this has led to the development of domain specific accelerators which enhance computing performance for specific applications [dally2020domain, jouppi2021tenlessons]. On the software level, attention has recently been brought to energy-aware algorithms for scientific computing. Mixed precision iterative refinement has successfully been applied in this area.
In [haidar2018design], the potential for using iterative refinement in designing energy-efficient linear solvers was studied in the context of the $LU$ factorisation of a dense matrix. The FP16-TC dhgesv-TC (Tensor Cores) solver using iterative refinement and half precision for the matrix factorisation achieved more than 5×5$\times$5 × improvement in terms of energy efficiency than then standard dgesv routine which factorises the matrix in double precision with no iterative refinement.
In [haidar2018harnessing], several numerical experiments measuring the performance gain of low precision arithmetic were conducted using an NVIDIA V100 PCIe GPU. The LAPACK algorithm GETRF for computing the $LU$ factorisation of a dense matrix was studied in various precisions. Their implementation computing the $LU$ factorisation accurately to half precision using TC achieves a speedup of 4×4$\times$4 × to 5×5$\times$5 × compared to double precision and a 2×2$\times$2 × speedup over single precision (both not using Tensor Cores).
In our experiments, we obtained a simulated gain of up to 3.5$\approx$ 3.5$\times$ 3.5 × in terms of memory references; see Figure 6. The ratio of the cost of a memory access to the cost of a floating point operation is continuing to grow in recent architectures [jouppi2021tenlessons, Table 2]. On modern 7nm semiconductor architectures, a cache (SRAM) memory access is on average 10×10$\times$10 × to 100×100$\times$100 × more expensive in terms of energy than a floating point operation. At the same time, accessing SRAM is up to 100×100$\times$100 × cheaper than accessing DRAM. Therefore, an efficient parallel implementation of iterative refinement within the proposed adaptive algorithm (Algorithm 4.1) that allows one to store more data in the cache memory can bring a significant reduction in energy cost, which is a promising area for future research.

### v3

# Exploiting Inexact Computations in Multilevel Monte Carlo and Other Sampling Methods

@@ Abstract
Multilevel sampling methods, such as multilevel and multifidelity Monte Carlo, multilevel stochastic collocation, or delayed acceptance Markov chain Monte Carlo, have become standard uncertainty quantification (UQ) tools for a wide class of forward and inverse problems. The underlying idea is to achieve faster convergence by leveraging a hierarchy of models, such as partial differential equation (PDE) or stochastic differential equation (SDE) discretisations with increasing accuracy. By optimally redistributing work among the levels, multilevel methods can achieve significant performance improvement compared to single level methods working with one high-fidelity model. Intuitively, approximate solutions on coarser levels can tolerate large computational error without affecting the overall accuracy. We show how this can be used in high-performance computing applications to obtain a significant performance gain.
As a use case, we analyse the computational error in the standard multilevel Monte Carlo method and formulate an adaptive algorithm which determines a minimum required computational accuracy on each level of discretisation. We show two examples of how the inexactness can be converted into actual gains using an elliptic PDE with lognormal random coefficients. Using a low precision sparse direct solver combined with iterative refinement results in a simulated gain in memory references of up to $3.5\times$ compared to the reference double precision solver; while using a MINRES iterative solver, a practical speedup of up to $1.5\times$ in terms of FLOPs is achieved. These results provide a step in the direction of energy-aware scientific computing, with significant potential for energy savings.
@@ 1 Introduction
Suppose we are interested in sampling from a probability distribution of a certain quantity $Q$, which depends on the (infinite-dimensional) solution of a partial differential equation (PDE) or a stochastic differential equation (SDE). In most cases, direct access to $Q$ is unavailable. Instead, a numerical model is used to obtain a finite-dimensional approximation $Q_{L}$ of the quantity $Q$. With increasing $L$ the accuracy of the model increases, but so does the cost of computing the solution. This can be a finite element or finite difference method in the case of PDEs, or an Euler-Maruyama discretisation for SDEs. A first idea to approximate the unavailable distribution of $Q$ is to choose a high-fidelity model $Q_{L}$ with $L$ sufficiently large and to use this model for sampling. Instead, the key idea of multilevel and multifidelity sampling methods, such as multilevel Monte Carlo (MLMC) [14], multilevel stochastic collocation [29], multilevel MCMC [11], or multi-index Monte Carlo (MIMC) [18], is to use a hierarchy of models $Q_{0},\ldots,Q_{L}$ with increasing accuracy. Then, by combining all the samples from models $Q_{0},\ldots,Q_{L}$ in a suitable way, a significant cost gain can be achieved. This paper adds one additional layer to the standard analysis, which has, to our knowledge, largely been neglected. In practice, the finite-dimensional model $Q_{l}$ is never solved exactly, rather an approximation $\tilde{Q}_{l}$ is obtained on a computer. We show that multilevel methods often admit safe use of inexact computations on coarse levels $l$ without affecting overall sampling accuracy. By employing techniques of energy-efficient computing this can lead to significant performance gains.
1.1 Inexact computations and energy-efficient computing
In the past decades, power has become the principal constraint for computing performance [24]. On the hardware level, this has led to the development of domain specific accelerators which enhance computing performance for specific applications [10, 25]. On the software level, attention has recently been brought to energy-aware algorithms for scientific computing. In recent years, there has been growing interest in the selective use of low precision floating point arithmetic to accelerate scientific computations while maintaining acceptable levels of accuracy; cf. [21].
An example of such an approach is mixed-precision iterative refinement, which can be used when the computation of the inexact $\tilde{Q}_{l}$ involves the solution of a linear system [20, 31]. The potential for using iterative refinement in designing energy-efficient linear solvers was studied in [17] in the context of $LU$ factorisation of a dense matrix. The FP16-TC dhgesv-TC (Tensor Cores) solver using iterative refinement and half precision for the matrix factorisation achieved more than $5\times$ improvement in terms of energy efficiency than the standard dgesv routine which factorises the matrix in double precision with no iterative refinement.
In principle, any iterative procedure can be used to lower the accuracy of the inexact solution $\tilde{Q}_{l}$. The optimal number of iterations can then be determined via a suitable stopping criterion – an approach independent of the iterative procedure. To demonstrate the flexibility of such an approach, we consider in this work both direct and iterative solvers for linear systems. To exploit inexact computations efficiently within direct methods, we employ mixed-precision iterative refinement.
We discuss in detail how these techniques can be used to improve the performance of multilevel sampling methods. We carry out a detailed analysis for multilevel Monte Carlo and comment on how a similar analysis can be done for multilevel Markov chain Monte Carlo [11] and multi-index Monte Carlo [18].
1.2 Case study: multilevel Monte Carlo
For a detailed numerical analysis we restrict ourselves to forward uncertainty quantification (UQ) and to multilevel Monte Carlo, tying the required computational accuracy to the discretisation error at level $l$. We choose a theoretical framework and a suitable error model to quantify the computational error. As a model problem, we consider an elliptic PDE with lognormal random coefficients of the form
$\nabla\cdot\bigl(a(\cdot,\omega)\nabla u(\cdot,\omega)\bigl)=f(\cdot,\omega)$
depending on the random parameter $\omega$. Such problems arise, for example, in UQ of groundwater flow [8]. The coefficient $a$ and the right-hand side $f$ are assumed to be (infinite-dimensional) random fields. Importantly, the dominant cost lies in the solution of a large system of linear algebraic equations.
Given a (scalar-valued) function $G$ of the solution $u$, such as the solution at a certain point in the domain, we are interested in sampling from the unavailable distribution of the quantity $Q=G(\omega)$ to compute statistics of this distribution, e.g., the mean $\mathbb{E}[Q]$. In practice, to obtain a computable approximation $Q_{L}$ of $Q$ we choose the finite element method (FEM), and to approximate the expected value $\mathbb{E}[Q_{L}]$ we employ a Monte Carlo (MC) method. This introduces a discretisation error and a sampling error. A significant variance reduction can be achieved if the samples are taken on a hierarchy of discretisation levels. This is the underlying idea of the MLMC method [19, 8, 14].
In our model problem, the dominant cost on each level of this hierarchy is the solution of the resulting FE system for each parameter $\omega$. The cost of sampling the input random fields $a(\cdot,\omega)$ and $f(\cdot,\omega)$ can be largely neglected. The error introduced by solving the linear system will be referred to as the computational error. The aim is to balance this computational error with the sampling and discretisation errors, and to employ techniques of energy-efficient computing to obtain performance gains.
This is in contrast to applications in computational finance, such as [15], which is concerned with the analysis of rounding errors in generating random variables in the context of stochastic differential equations (SDEs) and applications within MLMC. See also the earlier paper [4], where the authors explored the use of lower precision on field programmable gate arrays (FPGA) in the MLMC method for SDEs.
1.3 Main contributions and outline of the paper
In this paper, we
@@ 1.3 Main contributions and outline of the paper
  • propose a novel adaptive algorithm, which determines the minimum required computational accuracy on each level of discretisation using a-priori error estimates with no additional cost (Section 4.3);
  • provide a theoretical basis for applying this adaptive algorithm to an elliptic PDE with random coefficients and random right-hand sides (Section 4.5);
  • demonstrate the efficiency of this adaptive algorithm in a sequence of numerical examples, achieving up to a factor of about $3.5\times$ in simulated memory gain for iterative refinement and a speedup by a factor of about $1.5\times$ in terms of floating point operations in an iterative solver. A cost analysis and possible use in energy-efficient scientific computing are presented in Section 4.4.
The manuscript is divided into five sections. In Section 2, we discuss linear solvers and various types of errors they incur. We give an overview of floating point arithmetic, the technique of iterative refinement, as well as some convergence results for Krylov subspace methods. Section 3 introduces the elliptic model problem and its numerical solution via FEM. Selected FE convergence results are presented and the convergence of the FEM with inexact solvers is discussed. After a brief overview of the standard MLMC method, the new theoretical results are presented in Section 4, where we analyse the computational error in MLMC along with the adaptive algorithm and its computational complexity. Finally, the numerical results, as well as a thorough discussion are given in Section 5.
2 Inexact computations in linear solvers
@@ 2.1 Floating point arithmetic
A floating point (FP) number system $\mathbb{F}$ is a finite subset of real numbers whose elements can be written in a specific form; see [20] for a thorough description. Here, all computations in FP arithmetic are assumed to be carried out under the following standard model: For all $x,y\in\mathbb{F}$
@@
where $\varepsilon$ is the unit roundoff. Since most computations can be decomposed in terms of these basic operations, the above assumption allows us to analyse the error of a given algorithm. For simplicity, we will often abbreviate “the floating point arithmetic format with unit roundoff $\varepsilon$” to “the precision $\varepsilon$”.
The standard model is valid in particular for IEEE arithmetic, a technical standard of floating point arithmetic, which assumes the preliminaries above and adds other technical assumptions; see [20] for an overview. The IEEE standard defines several basic formats. In this work, we use formats both standardized and not standardised by IEEE. All of the formats we will use are hardware-supported, for instance by the NVIDIA H100 SXM GPU (for specifications see [27]). To be specific, in this manuscript we use quarter (q43), half, single, and double precision. Half, single, and double are basic IEEE formats. The quarter precision format we use has $4$ exponent bits and $3$ significand bits, which means storing one number requires $8$ bits including the sign. The unit roundoff of quarter precision is $2^{-4}$ and its range is $10^{\pm 2}$.
@@ 2.2 Krylov subspace methods
Krylov subspace methods are a powerful class of iterative methods for large-scale linear systems of equations and eigenvalue problems, particularly those involving sparse or structured matrices. Our application yields matrices which are symmetric positive definite (see Section 3), which makes the conjugate gradient (CG) and MINRES methods two suitable Krylov subspace methods for our purpose. The following overview is adapted from [16, Section 3.1].
Consider a linear system $Ax=b$ with symmetric positive definite $A\in\mathbb{R}^{n\times n}$. Let $x_{0}\in\mathbb{R}^{n}$ be an inital estimate of the solution, $e_{0}=x-x_{0}$ the initial error, and $r_{0}=b-Ax_{0}$ the initial residual. Then, in the $k$th iteration, CG and MINRES minimize the $A$-norm of the error and the Euclidean norm of the residual over the Krylov subspace $x_{0}+\text{span}\{r_{0},A^{2}r_{0},\dots,A^{k-1}r_{0}\}$, respectively.
The stopping criterion we use is on the relative residual. It is motivated by our specific use of the linear solver and will allow us to apply abstract error estimates from Section 4.2; see also Section 5. To be precise, we require the solution produced by the iterative algorithm to satisfy (1) for a given $\delta_{e}>0$, which motivates the choice of MINRES. It has been shown that for Hermitian matrices the residuals produced by CG and MINRES are closely related (see [16, Exercise 5.1]). Therefore, both methods are expected to perform similarly for our model problem in Section 3.
It is clear that in exact arithmetic, MINRES converges to the true solution in a finite number of iterations. In finite precision arithmetic, this is not the case, and the convergence analysis is nontrivial; see for example [16, Chapter 4]. We therefore perform all calculations in the MINRES method in double precision. In order to apply the abstract error estimates from Section 4.2 to MINRES, we assume $\widehat{x}_{k}\approx x_{k}$, where $\widehat{x}_{k}$ and $x_{k}$ are the MINRES solutions from $k$-th iteration computed in floating point and exact arithmetic, respectively. An alternative approach to performing all calculations in high precision would be to employ a Krylov subspace method coupled with iterative refinement; see Section 2.3 for an overview of iterative refinement. Iterative refinement as a technique to improve accuracy of a Krylov subspace method was used, for example, in [5] and [6] with the GMRES method.
@@ 2.3 Iterative refinement
Consider a linear system $Ax=b$ where $A\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$. Iterative refinement is a technique to enhance the accuracy of a numerical solution to this linear system by computing a residual vector $r$ and then correcting the current approximation by solving a linear system with right hand side $r$ to reduce the error; see [20, Section 12] for a detailed overview. In this way, the process can be repeated iteratively until a limiting level of accuracy is achieved. The particular version of iterative refinement used in this work is presented in Algorithm 2.1. It is a special case of [6, Algorithm 1.1].
@@
[…]
@@
1: Factorise AA in precision $\varepsilon_{f}$.
2: Solve $Ax_{0}=b$ in precision $\varepsilon_{f}$ by substitution and store $x_{0}$ in precision $\varepsilon$.
@@
4:  Compute $r_{i}=b-Ax_{i}$ in precision $\varepsilon_{r}$ and round $r_{i}$ to precision $\varepsilon_{s}$.
@@
8:  Solve $Ad_{i}=r_{i}$ by (forward/backward) substitution in precision $\varepsilon_{f}$ or $\varepsilon$(using the factorisation computed in step 1) and store $d_{i}$ in precision $\varepsilon$.
9:  $x_{i+1}=x_{i}+d_{i}$ in precision $\varepsilon$.
@@
Algorithm 2.1 uses the following three precisions:
  • $\varepsilon$ is the (working) precision that the data $A$, $b$ and solution $x$ are stored in,
  • $\varepsilon_{f}$ is the precision that the factorisation of $A$ is computed in,
  • $\varepsilon_{r}$ is the precision that residuals are computed in.
  • $\varepsilon_{s}$ is the precision that the correction equation is solved in. [^1] Note that in this subsection and in Section 3 the symbol $\varepsilon_{s}$ does not denote single precision here. We use it to keep the indices in accordance with [6], where the precisions are denoted by $_{r}, uu$, $u_{f}$, and $u_{s}$, respectively.
@@
The precisions $\varepsilon_{r}$, $\varepsilon$, $\varepsilon_{s}$, and $\varepsilon_{f}$ take only the values quarter, half, single, or double precision in this work, and we assume that $\varepsilon_{r}\leq\varepsilon\leq\varepsilon_{s}\leq\varepsilon_{f}$. Since we assume a direct linear solver with Cholesky factorisation, we effectively can choose $\varepsilon_{s}=\varepsilon_{f}$ in our case. From the perspective of the limiting tolerance for the residual norm $\|{r_{i}}\|/\|{b}\|$, no numerical benefit is obtained by choosing $\varepsilon_{s}<\varepsilon_{f}$. Nevertheless, it turns out that using a higher precision in step 8 can help in practice to reach the same desired tolerance in fewer iterations with negligible additional cost per iteration. There is also no reason to use extra precision to compute the residual in step 4 of the algorithm, since we are only interested in bounding the backward error. Computing the residual in higher precision is useful if bounding the forward error is of interest; see [6].
@@
Algorithm 2.1 is used extensively in this work to achieve the desired accuracy of the computed solution. Additional benefits arise in terms of cost. The key point is to reuse the factorisation computed in step 1 of the algorithm in step 8 of the algorithm. By reusing the factorisation, the accuracy of the solution can be improved with little additional cost, since the factorisation is typically the dominant part in terms of the required number of operations; see [31, Section 2.2].
@@
As in our Krylov subspace method, the stopping criterion in Algorithm 2.1 is on the relative residual norm. It is again motivated by our specific use of iterative refinement; see Sections 4.2 and 5. From the definition of the stopping criterion in step 5 it follows that if Algorithm 2.1 converges, then the produced solution is computed effectively to precision $\delta_{e}$ in the sense of (1). This will allow us to apply abstract error estimates from Section 4.2 to iterative refinement. The convergence of Algorithm 2.1 is analysed thoroughly in [6]. We will return to it in Section 3.2.